# AN ACCELERATED PRIMAL DUAL ALGORITHM WITH BACKTRACKING FOR DECENTRALIZED CONSTRAINED OPTIMIZATION

### v1

@@ 6. Convergence guarantees for D-APDB and the Proof of Main Results
which shows that $\mathbf{z}^{*}\triangleq(\mathbf{x}^{*},\theta^{*},\lambda^{*})\in\mathop{\bf dom}\phi\times\mathop{\bf dom}h\times\mathbb{R}^{n|\mathcal{E}|}$ is a saddle point of $\mathcal{L}$ as well. Since (6.28) holds for any saddle point $\hat{\mathbf{z}}$ of $\mathcal{L}$, we can replace $\hat{\mathbf{z}}$ with $\mathbf{z}^{*}$ and the result continues to hold. Accordingly, we have $\bar{a}_{k+1}\leq\bar{a}_{k}-b_{k}+c_{k}$ for $k\geq 0$ where $\{\bar{a}_{k}\}_{k\geq 0}$ is such that $\bar{a}_{k}=t_{k}Q^{k}(\mathbf{z}^{*})\geq 0$ for $k\geq 0$. Since $\sum_{k\geq 0}c_{k}<\infty$, we have $\bar{a}\triangleq\lim_{k\to\infty}\bar{a}_{k}$ exists. Moreover, the subsequence $\{k_{n}\}_{n\geq 0}$ defined above is independent of $\mathbf{z}^{*}$ or $\hat{\mathbf{z}}$; hence, it follows from the definition of $Q^{k}(\mathbf{z}^{*})$ in (4.6a) and (6.29) that $\lim_{n\to\infty}a_{k_{n}}=t_{k_{n}}Q^{k_{n}}(\mathbf{z}^{*})=0$. Since for a convergent sequence, every subsequence converges to the same limit point, we must have $\bar{a}=\lim_{k\to\infty}\bar{a}_{k}=0$. Thus, the inequality in (6.27) with $\hat{\mathbf{z}}$ replaced with $\mathbf{z}^{*}$ and taking the limit as $k\to\infty$ implies that $\lim_{k\to\infty}\mathbf{x}^{k}=\mathbf{x}^{*}$. Since we have already observed that $\mathbf{x}^{*}=\mathbf{1}_{N}\otimes x^{*}$ for some $x^{*}\in\mathbb{R}^{n}$, we have $\lim_{k\to\infty}x_{i}^{k}=x^{*}$ for all $i\in\mathcal{N}$. Furthermore, $t_{k}\geq\hat{t}>0$ for all $k\geq 0$ and $T_{K}\to\infty$ as $K\to\infty$ implies that the weighted sequence $\lim_{K\to\infty}\bar{x}_{i}^{K}=\lim_{K\to\infty}\bar{x}_{i}^{K}=x^{*}$; thus, $A\mathbf{x}^{*}=0$ implies that $\lim_{K\to\infty}\|A\bar{\mathbf{x}}^{K}\|=0$. Finally, taking the limit on both sides of (6.9) and (6.10) as $K\to\infty$ implies that $\varphi^{*}=\sum_{i\in\mathcal{N}}\varphi_{i}(x^{*})$ and $d_{-\mathcal{K}_{i}}\big(g_{i}(x^{*})\big)=0$, i.e., $g_{i}(x^{*})\in\mathcal{K}_{i}$, for all $i\in\mathcal{N}$, which shows that $x^{*}$ is an optimal solution to (1.1). On the other hand, setting $\theta_{i}=0$ within (6.35) implies that $\left\langle-g_{i}(x^{*}),~\theta_{i}^{*}\right\rangle\leq 0$ for $i\in\mathcal{N}$; moreover, $g_{i}(x^{*})\in\mathcal{K}_{i}$ and $\theta_{i}^{*}\in\mathcal{K}^{*}$ imply that $\left\langle-g_{i}(x^{*}),~\theta_{i}^{*}\right\rangle\geq 0$; thus, the complementary slackness conditions $\left\langle-g_{i}(x^{*}),~\theta_{i}^{*}\right\rangle=0$ hold for all $i\in\mathcal{N}$, which together with (6.36), $\theta_{i}^{*}\in\mathcal{K}_{i}^{*}$ and $g_{i}(x^{*})\in\mathcal{K}_{i}$ for all $i\in\mathcal{N}$, and $A\mathbf{x}^{*}=0$ together imply that $\theta^{*}=[\theta_{i}^{*}]_{i\in\mathcal{N}}$ is an optimal dual solution. ∎
@@ 7. Experiments
In this section, we provide numerical experiments to illustrate the performance of D-APDB and D-APDB0. The former algorithm can handle node-specific functional constraints defined by $g_{i}(\cdot)$; whereas the latter one is tailored for problems that are either unconstrained or with simple constraint sets onto which projections are cheap to compute. We test D-APDB over $\ell_{1}$-norm-regularized quadratically constrained quadratic programming (QCQP) problems while D-APDB0 is tested over $\ell_{1}$-norm-regularized unconstrained quadratic programming problems. For both cases, we test our algorithms on a random small-world network $\mathcal{G}=(\mathcal{N},\mathcal{E})$, i.e. we choose $|\mathcal{N}|$ edges creating a random cycle over nodes, then the remaining $|\mathcal{E}|-|\mathcal{N}|$ edges are selected uniformly at random. In our experiments, we set $|\mathcal{E}|=24$ edges and $|\mathcal{N}|=12$ nodes. Our experiments are run on a MacBook Air equipped with an Apple M2 CPU and 8 GB of unified memory. Our code is available at https://github.com/Qiushui-Xu/D-APDB.
@@
For this class of randomly generated QCQP problems, we compare our D-APDB with D-APD in terms of log relative suboptimality, relative consensus error and infeasibility of the average iterate sequence $\{\bar{x}^{k}\}_{k\geq 0}\subset\mathbb{R}^{n}$, i.e., in Fig. 1, we plot $\log(|\varphi(\bar{x}^{k})-\varphi^{*}|/|\varphi^{*}|+1)$ in (A), $\sum_{i\in\mathcal{N}}\|x_{i}^{k}-\bar{x}^{k}\|^{2}/(N\|\bar{x}^{k}\|^{2})$ in (B) and $\max_{i\in\mathcal{N}}\|(g_{i}(\bar{x}^{k}))_{+}\|/\max_{i\in\mathcal{N}}\|(g_{i}(\bar{x}^{0}))_{+}\|$ in (C), where $\bar{x}^{k}=\sum_{i\in\mathcal{N}}x_{i}^{k}/N\in\mathbb{R}^{n}$ for $k\geq 0$. The benchmark algorithm D-APD is essentially our proposed method D-APDB without backtracking; that is, it uses a constant stepsize $\bar{\tau}_{i}$ and sets $\eta_{i}^{k}=1$ for all $k\geq 0$, rather than performing backtracking on these parameters. [^5] We note that D-APD is closely related to the method proposed in [hamedani2021decentralized] for decentralized constrained strongly convex problems when the strong convexity parameter is set to zero. Comparing against this benchmark allows us to assess the benefits of incorporating backtracking.
For the benchmark algorithm D-APD which is based on constant stepsize, we set the initial stepsizes $\bar{\tau}_{i}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\tau}_{i}}$ where $\hat{\tau}_{i}$ is as in eq. 5.7, $\zeta_{i}=1$, $c_{\alpha},c_{\beta},c_{\varsigma}=0.1$ and $\eta^{k}=1$. This stepsize requires knowledge of the Lipschitz constants $L_{f_{i}}$ and constants $L_{g_{i}}$; therefore, we compute these constants using the randomly generated the matrices $A_{i},Q_{i}$ [^6] Note $\nabla f_{i}(x)=Q_{i}x$; hence, we can take $L_{f_{i}}=\|Q_{i}\|_{2}$ and $L_{g_{i}}$ can be computed similarly as $L_{g_{i}}=\|A_{i}\|_{2}$.. For D-APDB, because it can potentially support larger stepsizes that adapts to the local curvature, we set the initial stepsizes to be larger than $\hat{\tau}_{i}$. More specifically, we set the initial stepsize of D-APDB (in Line 2 of D-APDB) as $\bar{\tau}_{i}=\kappa\hat{\tau}_{i}$ with $\kappa=20$, and we set the shrink factor $\rho=0.9$, $\zeta_{i}=1$, $c_{\alpha},c_{\beta},c_{\varsigma}=0.1$, and $\delta=0.1$. All the other parameters are set according to our theoretical results. The results in fig. 1 shows that our method outperforms D-APD on an average of 20 simulations, where the shaded areas demonstrate the standard deviation of the performance over the runs.
@@
Figure 1. Comparison of D-APDB  against D-APD for solving eq. 7.1 over 20 runs. “Average number of gradient calls per node” means $\frac{1}{N}\sum_{i=1}^{N}n_{i}^{k}$ where $n_{i}^{k}$ counts how many calls for gradients for node $i$ at iteration $k$.
@@ 7.2. Unconstrained $\ell_{1}$-norm-regularized quadratic programming
For this class of randomly generated unconstrained $\ell_{1}$-norm-regularized QP problems, we compare our D-APDB0 with D-APD and global_DATOS [chen2025parameter] in terms of log relative suboptimality, relative consensus error of the iterate sequence, i.e., in the two panels of Fig. 2, from left to right, we plot $\log(|\varphi(\bar{x}^{k})-\varphi^{*}|/|\varphi^{*}|+1)$, $\sum_{i=1}^{N}\|x_{i}^{k}-\bar{x}^{k}\|^{2}/(N\|\bar{x}^{k}\|^{2})$. Benchmarking against these methods allows us to evaluate the benefits of incorporating node-specific backtracking and employing node-specific stepsize choices. global_DATOS uses two communication rounds among the neighboring nodes per iteration iteration while both D-APDB0 and D-APD require one communication round per iteration. In Fig. 2 we plotted sub-optimality and consensus violation against the number of communication rounds in the x-axis.
In the experiments, for the benchmark algorithm D-APD, we set the initial stepsizes to $\hat{\tau}_{i}=\frac{1}{2L_{f_{i}}}$ for each agent $i$, and choose $c_{\alpha}=c_{\varsigma}=0.4$ and $\eta_{i}^{k}=1$ for all $k$. For D-APDB0, because it can support larger stepsizes that adapt to the local curvature, we use a more aggressive initialization. Specifically, in Line 2 of D-APDB0, we set the initial stepsizes to $\bar{\tau}_{i}=\kappa\hat{\tau}_{i}$ with $\kappa=5$, and choose the shrink factor $\rho=0.9$, $c_{\alpha}=c_{\varsigma}=0.4$, and $\delta=0.1$. For global_DATOS, node-specific stepsizes are not allowed, but backtracking enables the use of larger global stepsizes. Therefore, we set the initial stepsize for all nodes to $\bar{\alpha}=\kappa\hat{\alpha}$, where $\hat{\alpha}\triangleq\max_{i\in\mathcal{N}}\hat{\tau}_{i}$ and again choose $\kappa=5$ for a fair comparison between D-APDB0 and global_DATOS. We follow the definition of Laplacian-based constant edge weight matrix $W=I-\frac{\Omega}{d_{\rm max}+1}$ in [shi2015extra] to generate the gossip matrix for global_DATOS, where $\Omega$ is the Laplacian matrix of the graph $\mathcal{G}$ and $d_{\rm max}$ is the max degree of the graph $\mathcal{G}$. All other parameters follow the theoretical prescriptions. The results in fig. 2 show that our method consistently outperforms both D-APD and global_DATOS over an average of 20 simulations, where the shaded regions indicate the standard deviation across runs.
@@
Figure 2. Comparison of D-APDB0  against D-APD and global_DATOS for solving eq. 7.2 with 20 times simulation.

### v2

@@ 6. Convergence guarantees for D-APDB and the Proof of Main Results
which shows that $\mathbf{z}^{*}\triangleq(\mathbf{x}^{*},\theta^{*},\lambda^{*})\in\mathop{\bf dom}\phi\times\mathop{\bf dom}h\times\mathbb{R}^{n|\mathcal{E}|}$ is a saddle point of $\mathcal{L}$ as well. Since (6.4) holds for any saddle point $\hat{\mathbf{z}}$ of $\mathcal{L}$, we can replace $\hat{\mathbf{z}}$ with $\mathbf{z}^{*}$ and the same arguments we used in the proof of Theorem 6.4 continue to hold. Accordingly, we have $\bar{a}_{k+1}\leq\bar{a}_{k}-b_{k}+c_{k}$ for $k\geq 0$ where $\{\bar{a}_{k}\}_{k\geq 0}$ is such that $\bar{a}_{k}=t_{k}Q^{k}(\mathbf{z}^{*})\geq 0$ for $k\geq 0$. Since $\sum_{k\geq 0}c_{k}<\infty$, we have $\bar{a}\triangleq\lim_{k\to\infty}\bar{a}_{k}$ exists. Moreover, the subsequence $\{k_{n}\}_{n\geq 0}$ defined above is independent of $\mathbf{z}^{*}$ or $\hat{\mathbf{z}}$; hence, it follows from the definition of $Q^{k}(\mathbf{z}^{*})$ in (4.5a) and (6.22) that $\lim_{n\to\infty}a_{k_{n}}=t_{k_{n}}Q^{k_{n}}(\mathbf{z}^{*})=0$. Since for a convergent sequence, every subsequence converges to the same limit point, we must have $\bar{a}=\lim_{k\to\infty}\bar{a}_{k}=0$. Thus, the inequality in (6.3) with $\hat{\mathbf{z}}$ replaced with $\mathbf{z}^{*}$ and taking the limit as $k\to\infty$ implies that $\lim_{k\to\infty}\mathbf{x}^{k}=\mathbf{x}^{*}$. Since we have already observed that $\mathbf{x}^{*}=\mathbf{1}_{N}\otimes x^{*}$ for some $x^{*}\in\mathbb{R}^{n}$, we have $\lim_{k\to\infty}x_{i}^{k}=x^{*}$ for all $i\in\mathcal{N}$. Furthermore, $t_{k}\geq\hat{t}>0$ for all $k\geq 0$ and $T_{K}\to\infty$ as $K\to\infty$ implies that the weighted sequence $\lim_{K\to\infty}\bar{x}_{i}^{K}=\lim_{K\to\infty}\bar{x}_{i}^{K}=x^{*}$; thus, $A\mathbf{x}^{*}=0$ implies that $\lim_{K\to\infty}\|A\bar{\mathbf{x}}^{K}\|=0$. Finally, taking the limit on both sides of (6.9) and (6.10) as $K\to\infty$ implies that $\varphi^{*}=\sum_{i\in\mathcal{N}}\varphi_{i}(x^{*})$ and $d_{-\mathcal{K}_{i}}\big(g_{i}(x^{*})\big)=0$, i.e., $g_{i}(x^{*})\in\mathcal{K}_{i}$, for all $i\in\mathcal{N}$, which shows that $x^{*}$ is an optimal solution to (1.1). On the other hand, setting $\theta_{i}=0$ within (6.28) implies that $\left\langle-g_{i}(x^{*}),~\theta_{i}^{*}\right\rangle\leq 0$ for $i\in\mathcal{N}$; moreover, $g_{i}(x^{*})\in\mathcal{K}_{i}$ and $\theta_{i}^{*}\in\mathcal{K}^{*}$ imply that $\left\langle-g_{i}(x^{*}),~\theta_{i}^{*}\right\rangle\geq 0$; thus, the complementary slackness conditions $\left\langle-g_{i}(x^{*}),~\theta_{i}^{*}\right\rangle=0$ hold for all $i\in\mathcal{N}$, which together with (6.29), $\theta_{i}^{*}\in\mathcal{K}_{i}^{*}$ and $g_{i}(x^{*})\in\mathcal{K}_{i}$ for all $i\in\mathcal{N}$, and $A\mathbf{x}^{*}=0$ together imply that $\theta^{*}=[\theta_{i}^{*}]_{i\in\mathcal{N}}$ is an optimal dual solution. ∎
@@ 7. Experiments
In this section, we provide numerical experiments to illustrate the performance of D-APDB and D-APDB0. The former algorithm can handle node-specific functional constraints defined by $g_{i}(\cdot)$; whereas the latter one is tailored for problems that are either unconstrained or with simple constraint sets onto which projections are cheap to compute. We test D-APDB over $\ell_{1}$-norm-regularized quadratically constrained quadratic programming (QCQP) problems and on primal support vector machine (SVM) training problems while D-APDB0 is tested over $\ell_{1}$-norm-regularized unconstrained quadratic programming problems. For all experiments, we test our algorithms on a random small-world network $\mathcal{G}=(\mathcal{N},\mathcal{E})$, i.e., we choose $|\mathcal{N}|$ edges creating a random cycle over nodes, then the remaining $|\mathcal{E}|-|\mathcal{N}|$ edges are selected uniformly at random. In our experiments, we set $|\mathcal{E}|=24$ edges and $|\mathcal{N}|=12$ nodes. Our experiments are run on a MacBook Air equipped with an Apple M2 CPU and 8 GB of unified memory. Our code is available at https://github.com/Qiushui-Xu/D-APDB.
@@
For this class of randomly generated QCQP problems, we compare our D-APDB with D-APD in terms of log relative suboptimality, relative consensus error and infeasibility of the average iterate sequence $\{\bar{x}^{k}\}_{k\geq 0}\subset\mathbb{R}^{n}$, i.e., in Fig. 1, we plot $\log(|\varphi(\bar{x}^{k})-\varphi^{*}|/|\varphi^{*}|+1)$ in (A), $\sum_{i\in\mathcal{N}}\|x_{i}^{k}-\bar{x}^{k}\|^{2}/(N\|\bar{x}^{k}\|^{2})$ in (B) and $\max_{i\in\mathcal{N}}\|(g_{i}(\bar{x}^{k}))_{+}\|/\max_{i\in\mathcal{N}}\|(g_{i}(\bar{x}^{0}))_{+}\|$ in (C), where $\bar{x}^{k}=\sum_{i\in\mathcal{N}}x_{i}^{k}/N\in\mathbb{R}^{n}$ for $k\geq 0$. The benchmark algorithm D-APD is essentially our proposed method D-APDB without backtracking; that is, it uses a constant stepsize $\bar{\tau}_{i}$ and sets $\eta_{i}^{k}=1$ for all $k\geq 0$, rather than performing backtracking on these parameters5.5We note that D-APD is closely related to the method proposed in [19] for decentralized constrained strongly convex problems when the strong convexity parameter is set to zero. Comparing against this benchmark allows us to assess the benefits of incorporating backtracking.
For the benchmark algorithm D-APD which is based on constant stepsize, we set the initial stepsizes $\bar{\tau}_{i}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\tau}_{i}}$ where $\hat{\tau}_{i}$ is as in (5.7), $\zeta_{i}=1$, $c_{\alpha},c_{\beta},c_{\varsigma}=0.1$ and $\eta^{k}=1$ for $k\geq 0$. This stepsize requires knowledge of the Lipschitz constants $L_{f_{i}}$ and constants $L_{g_{i}}$; therefore, we compute these constants using the randomly generated matrices $A_{i},Q_{i}$66Note $\nabla f_{i}(x)=Q_{i}x$; hence, we can take $L_{f_{i}}=\|Q_{i}\|_{2}$ and $L_{g_{i}}$ can be computed similarly as $L_{g_{i}}=\|A_{i}\|_{2}$.. For D-APDB, because it can potentially support larger stepsizes that adapt to the local curvature, we set the initial stepsizes to be larger than $\hat{\tau}_{i}$. More specifically, we set the initial stepsize of D-APDB (in Line 2 of D-APDB) as $\bar{\tau}_{i}=\kappa\hat{\tau}_{i}$ with $\kappa=20$, and we set the shrink factor $\rho=0.9$, $\zeta_{i}=1$, $c_{\alpha},c_{\beta},c_{\varsigma}=0.1$, and $\delta=0.1$. All the other parameters are set according to our theoretical results. The results in Fig. 1 shows that our method outperforms D-APD on average over 20 simulations, each corresponding to a randomly generated problem instance, where the shaded areas demonstrate the standard deviation of the performance over the runs.
@@
Figure 1. Comparison of D-APDB  against D-APD for solving (7.1) over 20 runs. “Average number of gradient calls per node” means $\frac{1}{N}\sum_{i=1}^{N}n_{i}^{k}$ where $n_{i}^{k}$ counts how many calls for gradients for node $i$ at iteration $k$.
@@ 7.2. Unconstrained $\ell_{1}$-norm-regularized quadratic programming
For this class of randomly generated unconstrained $\ell_{1}$-norm-regularized QP problems, we compare our D-APDB0 with D-APD and global_DATOS [14] in terms of log relative suboptimality, relative consensus error of the iterate sequence, i.e., in the two panels of Fig. 2, from left to right, we plot $\log(|\varphi(\bar{x}^{k})-\varphi^{*}|/|\varphi^{*}|+1)$, $\sum_{i=1}^{N}\|x_{i}^{k}-\bar{x}^{k}\|^{2}/(N\|\bar{x}^{k}\|^{2})$. Benchmarking against these methods allows us to evaluate the benefits of incorporating node-specific backtracking and employing node-specific stepsize choices. global_DATOS uses two communication rounds among the neighboring nodes per iteration while both D-APDB0 and D-APD require one communication round per iteration. In Fig. 2 we plotted sub-optimality and consensus violation against the number of communication rounds in the x-axis.
In the experiments, for the benchmark algorithm D-APD, we set the initial stepsizes to $\bar{\tau}_{i}=\frac{1}{2L_{f_{i}}}$ for each $i\in\mathcal{N}$, and choose $c_{\alpha}=c_{\varsigma}=0.4$ and $\eta_{i}^{k}=1$ for all $k\geq 0$. For D-APDB0, because it can support larger stepsizes, we use a more aggressive initialization. Specifically, in Line 2 of D-APDB0, we set the initial stepsizes to $\bar{\tau}_{i}=\kappa\hat{\tau}_{i}$ with $\kappa=5$, and choose the shrink factor $\rho=0.9$, $c_{\alpha}=c_{\varsigma}=0.4$, and $\delta=0.1$. For global_DATOS, node-specific stepsizes are not allowed, but backtracking enables the use of larger global stepsizes. Therefore, we set the initial stepsize for all nodes to $\bar{\alpha}=\kappa\hat{\alpha}$, where $\hat{\alpha}\triangleq\max_{i\in\mathcal{N}}\hat{\tau}_{i}$ and again choose $\kappa=5$ for a fair comparison between D-APDB0 and global_DATOS. We follow the definition of Laplacian-based constant edge weight matrix $W=I-\frac{\Omega}{d_{\rm max}+1}$ in [49] to generate the gossip matrix for global_DATOS, where $\Omega$ is the Laplacian matrix of the graph $\mathcal{G}$ and $d_{\rm max}$ is the max degree of the graph $\mathcal{G}$. All other parameters follow the theoretical prescriptions. The results in Fig. 2 show that our method consistently outperforms both D-APD and global_DATOS on average for 20 simulations, where the shaded regions indicate the standard deviation across runs.
@@
Figure 2. Comparison of D-APDB0  against D-APD and global_DATOS for solving (7.2) with 20 simulation.
(a) Log Relative Suboptimality
(b) Consensus Error
(c) Average Constraint Violation
(d) Log Relative Suboptimality
(e) Consensus Error
(f) Average Constraint Violation
Figure 3. Comparison of D-APDB  against D-APD for solving the linear SVM problem in (7.3) over 20 random problem instances. “Average number of gradient calls per node” means $\frac{1}{N}\sum_{i=1}^{N}n_{i}^{k}$ where $n_{i}^{k}$ counts how many calls for gradients for node $i$ at iteration $k$.
7.3. Primal linear SVM problems.
We consider a primal linear SVM problem with data distributed among computing nodes in $\mathcal{N}$. Let $\mathcal{S}\triangleq\{1,2,\cdots,s\}$ denote the set of indices for training data and $\mathcal{D}\triangleq\{(x_{\ell},y_{\ell})\in\mathbb{R}^{n}\times\{-1,+1\}:\ell\in\mathcal{S}\}$ be a set of feature vector and label pairs. Let $\{\mathcal{S}_{i}\}_{i\in\mathcal{N}}$ be a partition of $\mathcal{S}$ among the nodes $\mathcal{N}$.
Define $\mathbf{w}=[w_{i}]_{i\in\mathcal{N}}$, $\mathbf{b}=[b_{i}]_{i\in\mathcal{N}}$, and $\varepsilon\in\mathbb{R}^{s}$ such that $w_{i}\in\mathbb{R}^{n}$ and $b_{i}\in\mathbb{R}$ for $i\in\mathcal{N}$. Consider the following distributed SVM problem:
$\varphi(\mathbf{w},\mathbf{b})\triangleq\min_{\mathbf{w},\mathbf{b},\varepsilon}$ $\ \frac{1}{2}\sum_{i\in\mathcal{N}}\|w_{i}\|^{2}+NC\sum_{i\in\mathcal{N}}\sum_{\ell\in\mathcal{S}_{i}}\varepsilon_{\ell}$ (7.3)
${\rm s.t.}$ $\quad y_{\ell}(w_{i}^{\top}x_{\ell}+b_{i})\geq 1-\varepsilon_{\ell},\quad\varepsilon_{\ell}\geq 0,\quad\ell\in\mathcal{S}_{i},\quad i\in\mathcal{N},$
$\quad w_{i}=w_{j},\quad b_{i}=b_{j},\quad(i,j)\in\mathcal{E}.$
Similar to [17], $\{x_{\ell}\}_{\ell\in\mathcal{S}}$ is generated from two-dimensional multivariate Gaussian distribution with covariance matrix $\Sigma=[1,0;0,2]$ and with mean vector either $m_{1}=[-1,-1]^{\top}$ or $m_{2}=[1,1]^{\top}$ with equal probability. The experiment was performed for $s=600$ training points such that for each $i\in\mathcal{N}$, we randomly sample $50$ training data points from $\mathcal{S}$, i.e., $|\mathcal{S}_{i}|=50$ for $i\in\mathcal{N}$. For this class of randomly generated primal linear SVM problems, we compare our D-APDB with D-APD in terms of log relative suboptimality, consensus error and average constraint violation of the iterate sequence, i.e., in Fig. 3, we plot $\log(|\varphi(\bar{w}^{k},\bar{b}^{k})-\varphi^{*}|/|\varphi^{*}|+1)$ in (A), $\frac{1}{N}\sum_{i\in\mathcal{N}}\|[w_{i}^{k},b_{i}^{k}]-[\bar{w}^{k},\bar{b}^{k}]\|^{2}$ in (B) and $\frac{1}{s}\sum_{i\in\mathcal{N}}\sum_{\ell\in\mathcal{S}_{i}}(1-\epsilon^{k}_{\ell}-y_{\ell}(\bar{w}^{k^{\top}}x_{\ell}+\bar{b}^{k}))_{+}$ in (C), where $\bar{w}^{k}=\sum_{i\in\mathcal{N}}w_{i}^{k}/N\in\mathbb{R}^{n}$ and $\bar{b}^{k}=\sum_{i\in\mathcal{N}}b_{i}^{k}/N\in\mathbb{R}$ for $k\geq 0$.
For the benchmark algorithm D-APD, which uses constant stepsize, we set the step size $\bar{\tau}_{i}=0.005$ (the first row in Fig. 3) to mimic the scenario where one picks a conservative stepsize in case the Lipschitz constants are not readily available. For the case when one knows the Lipschitz constants, we set $\bar{\tau}_{i}=\min\Big\{\frac{1-(\delta+c)}{2L_{f_{i}}},~\frac{1}{C_{g_{i}}}\sqrt{\frac{{c_{\alpha}(1-\delta)}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\zeta_{i}}}\Big\}$ shown in the second row of Fig. 3. For both cases we set $\zeta_{i}=1$ (for dual step size choice), $c_{\alpha},c_{\beta},c_{\varsigma}=0.1$ and $\eta^{k}=1$ for $k\geq 0$. For D-APDB, because it can potentially support larger stepsizes, we set the initial stepsizes to be larger than $\bar{\tau}_{i}$ choices for D-APD stated above. More specifically, we set the initial stepsize of D-APDB (in Line 2 of D-APDB) as 10 and 5 times larger than $\bar{\tau}_{i}$ of D-APD for the first and second scenario, respectively. Moreover, we set the contraction factor $\rho=0.9$, $\zeta_{i}=1$, $c_{\alpha},c_{\beta},c_{\varsigma}=0.1$, and $\delta=0.1$ for both settings. All the other parameters are set according to our theoretical results. The results in Fig. 3 shows that our method outperforms D-APD on average over 20 randomly generated problem instances, where the shaded areas demonstrate the standard deviation of the performance over the runs.